OpenStudy (anonymous):

**GIVING MEDAL** What is an equation of the line in slope intercept form? m = 4 and the y-intercept is (0, 5) y = 4x - 5 y = -5x + 4 y = 4x + 5 y = 5x + 4

OpenStudy (anonymous):

equation when a point is given and slope is, (y-y1)=m(x-x1) where m is the slope and x1,y1 is the given point

OpenStudy (anonymous):

(y-5)=4(x-0) y-5=4x y=4x+5